Directrix (directrix):

Fill in the numbers (exact value) that will make each of the six equations true.

Directrix (directrix):

Diagram attached.

OpenStudy (kdr1308):

I think --> \[\sqrt{18} - \sqrt{2} = 2\sqrt{2}\] I'm not 100% sure

jhonyy9 (jhonyy9):

sqrt18=3sqrt2 3sqrt2 -sqrt2 = 2sqrt2 2sqrt2 +sqrt2 /6 = 13sqrt2 /6 there is divided 2sqrt2 by sqrt2 /6 = ?

jhonyy9 (jhonyy9):

so than 2sqrt2 / sqrt2 /6 = 12sqrt2/sqrt2 =12

jhonyy9 (jhonyy9):

3sqrt2 /x = 54 3sqrt2 =54x 3sqrt2/54 = x sqrt2/18 = x

jhonyy9 (jhonyy9):

sqrt2/18 +x = sqrt2/6 x= sqrt2/6 -sqrt2/18 x= 2sqrt2 /18

jhonyy9 (jhonyy9):

sqrt2 *2sqrt2/18 = 4/18 = 2/9 54 2/9 = 12 54 / (2/9) = 12

jhonyy9 (jhonyy9):

@Directrix hope this will like you and that is right solved sure

Directrix (directrix):

I think so. @jhonyy9 @Kdr1308 Your first row is correct.